# OpticalAging: Real-time Presbyopia Simulation for Inclusive Design via Tunable Lenses

### v1

@@ Abstract.
Presbyopia, a common age-related vision condition affecting most people as they age, often remains inadequately understood by those unaffected. To help bridge the gap between abstract accessibility knowledge and a more grounded appreciation of perceptual challenges, this study presents OpticalAging, an optical see-through simulation approach. Unlike VR-based methods, OpticalAging uses dynamically controlled tunable lenses to simulate the first-person visual perspective of presbyopia’s distance-dependent blur during real-world interaction, aiming to enhance awareness. While acknowledging critiques regarding simulation’s limitations in fully capturing lived experience, we position this tool as a complement to user-centered methods. Our user study (N = 19, 18-35 years old) provides validation: quantitative measurements show statistically significant changes in near points across three age modes (40s, 50s, 60s), while qualitative results suggest increases in reported understanding and empathy among participants. The integration of our tool into a design task showcases its potential applicability within age-inclusive design workflows when used critically alongside direct user engagement.
Visual Simulation, Presbyopia, Visual Experience Modulation, Programmable Vision, Optical See-through, Accessibility, Inclusive Design, Aged Vision, Assistive Technology
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Accessibility systems and toolsccs: Human-centered computing Accessibility design and evaluation methodsccs: Human-centered computing Systems and tools for interaction design
@@ 1. Introduction
Presbyopia, an age-related visual impairment, reduces the ability to focus on near objects (BMC, 2021). It is a condition that virtually no one escapes (Polat et al., 2012; Charman, 2017; Wolffsohn et al., 2020; Heys et al., 2007; Atchison, 2008). Although traditional corrections such as spectacles or contact lenses offer functional solutions, they can introduce their own challenges (Katz et al., 2021; Wolffsohn et al., 2023). For instance, constantly needing to switch between different pairs of glasses can be cumbersome, multifocal lenses often require an adaptation period and may introduce peripheral distortion, and contact lenses might be uncomfortable or impractical for certain activities or environments. This impact underscores a critical gap in inclusive design: often stems from a lack of deep understanding among those unaffected by presbyopia (Sprung et al., 2022; Singer and Tusche, 2014). Addressing this understanding gap necessitates approaches that can help non-presbyopic individuals, particularly designers, appreciate the perceptual challenges associated with presbyopia, fostering broader awareness (Goertz et al., 2014). Bridging this gap requires tools that move beyond abstract design guidelines, helping designers develop a more grounded perspective than abstract knowledge alone might provide.
Advances in visual simulation technologies, including virtual reality (VR) and augmented reality (AR), have opened new avenues for simulating visual impairments. VR has shown versatility in simulating conditions and enhancing understanding (VRH, 2021; Alexander et al., 2020). However, while VR offers versatile simulation platforms, its inherent disconnect from the physical world limits its utility for crucial design evaluation tasks that involve interacting with physical prototypes, assessing ergonomics directly, or judging legibility on actual product surfaces and packaging within real-world environments. Optical see-through approaches, in contrast, allow simulation directly within the user’s real-world context, enabling designers to experience simulated impairments while handling tangible mock-ups and interacting naturally with the intended use environment (Zhang et al., 2022; Krösl et al., 2020; Padmanaban et al., 2018; Langlotz et al., 2018).
@@
To validate our approach, we conducted a user study with three objectives: (1) quantitatively examine simulated presbyopia in three age modes (40s, 50s, and 60s); (2) assess the reported change in understanding and empathy following interaction with the simulation; and (3) investigate the practical application of our simulation in real-world design scenarios with a professional designer and a design student. The results demonstrate that our approach significantly alters the near points of the users, enabling them to experience specific visual effects associated with presbyopia. Our system provides a perceptually relevant simulation of distance-dependent blur. Participants reported observations about daily challenges associated with these visual effects, suggesting increases in reported empathy and understanding. An exploratory case study with designer participants revealed the potential of our system to inform aspects of design processes, contributing towards more intuitive, user-centered solutions for aging populations.
This research contributes to the field of inclusive design support in several ways: (1) A novel optical see-through system using tunable lenses for first-person, real-time simulation of specific visual effects associated with presbyopia, enabling direct evaluation of real-world interactions. (2) A method intended to help bridge the gap between abstract accessibility guidelines and a more grounded appreciation of perceptual challenges relevant to inclusive design for aging populations. (3) Strong quantitative validation of the simulation’s ability to accurately shift near points, complemented by qualitative evidence suggesting potential enhancements in reported understanding and empathy among non-presbyopic users when interpreted within the limitations of simulation. (4) Explores the practical applications of presbyopia simulation in real-world design scenarios, offering insight into how such tools might be integrated into professional design workflows when used critically and supplementally.
@@ 2.1. Presbyopia: Background and Treatments
2.2. Visual Impairment Simulation
To improve understanding and insight into others’ visual conditions, various methods for simulating these conditions are explored. Techniques include methods based on static images (ZEI, EISS; honolulueyeclinic, 2024; Asada, 2011), interactive web demonstrations (Versanthealth, 2024), and commercial software products (2EyesVision, 2019; Media, 2022). Simpler optical methods, such as wearing goggles (Glasses, 2025; Design, 2025) with fixed diffusing lenses or static filters, have also been used to approximate certain visual impairments, such as general blur. However, such non-dynamic approaches cannot adequately simulate conditions like presbyopia, where the visual effect (blur) is fundamentally dependent on the viewing distance. Capturing the dynamic, distance-dependent nature of presbyopic blur and the associated difficulty in shifting focus requires a real-time, adjustable optical system.
Head-mounted displays (HMDs) are also effective tools for simulating visual conditions. In academic research, Jones et al. (Jones et al., 2020) investigated the use of virtual and augmented reality (VR/AR) to replicate common visual impairment symptoms, such as field loss, color reduction, glare, and diminished visual acuity. Additionally, Alexander et al. (Alexander et al., 2020) recommended a flexible approach using eye tracking to imitate usual symptoms of visual impairment. To achieve more realistic and intuitive simulations of visual impairments, Krosl et al. (Krösl et al., 2020) incorporated eye tracking and cataract simulations using pre-recorded 360-degree videos in HMD. While these studies have highlighted the significance of visual effects linked to eye movements and their implications in real-world environments. Zhang et al. (Zhang et al., 2022) advanced this work by employing optical see-through eyewear to replicate age-related macular degeneration and open-angle glaucoma, offering deeper insights into visually impaired individuals’ perspectives. They emphasized the immediate adaptiveness of optical see-through eyewear for simulating visual impairments. In the latest research, Orlosky et al. (Orlosky et al., 2024) proposed a distance-dependent blurring technique in virtual environments. Despite acknowledging past contributions to blurred visual effects, we note that the distance-dependent aspect of presbyopia remains insufficiently studied, particularly requiring highly realistic optical see-through applications.
While various simulation techniques exist, their application in design practice is complex and warrants careful consideration. Empirical research indicates that designers often utilize simulations, including those for visual impairments, due to practical workplace constraints such as limited time, budget, and significant challenges in recruiting or accessing disabled participants for direct feedback or co-design (Tigwell, 2021). Furthermore, studies suggest that many designers may not be fully aware of the significant critiques and potential pitfalls of simulation methods, such as oversimplification, the inability to capture lived expertise or compensatory strategies, and the risk of reinforcing stereotypes (Tigwell, 2021; Bennett and Rosner, 2019). Understanding this context – the practical pressures faced by designers alongside the valid concerns regarding simulation efficacy and ethics – is crucial when considering the development and deployment of new simulation tools like the one proposed in this paper.
2.3. Tunable Lenses for Vision Modulation
Mompeán et al. (Mompeán et al., 2020) and Padmanabhan et al. (Padmanaban et al., 2018) utilized tunable lenses to demonstrate the tuning of focus for correcting presbyopia. This work is particularly relevant to our research as it provides practical insights into applying tunable lenses for presbyopia correction. Most recently and similarly, Agarwala et al. (Agarwala et al., 2022) assessed electronically tunable lenses for their aberration properties and use as correction lenses. While these advancements focus on compensating for presbyopia in individuals who experience it, a distinct need exists to foster deeper understanding and empathy among those unaffected, particularly designers and developers creating solutions for aging populations. Correction technologies address the functional impairment, but simulation tools like ours aim to bridge the experiential gap, offering a first-person perspective intended to inform more inclusive and user-centered design practices. However, to our knowledge, no research optically simulated presbyopia in first-person-view for this purpose, as we explored in this research in the following sections.
@@ 3. Simulating Presbyopia through Dynamic Adjustment of Diopters using Tunable Lenses
Specification category Lens dimensions (Ø x thickness) Range of optical power Clear aperture Operating temperature Pupillary distance
Specification detail 15.8mm x 7.4mm -15D $\sim$ 15D 6.3mm 0 $\sim$ 45°C 55mm $\sim$ 70mm
Table 1. Tunable Lenses (built-in ViXion01 smart eyewear) Specifications.
@@
Algorithm: As presbyopia gradually progresses with age and individuals typically begin to notice symptoms in their 40s (Duane, 1922; Presbyopia, 2016), our simulation modes commence at this age, extending through the 50s and 60s, in line with previous research (Duane, 1912; Schwartz, 2013). To simulate presbyopic visual blur, the adjustable lenses alter the cumulative focal power of the wearer. The algorithm operates simply: If the tunable lens eyewear’s distance sensor identifies that the viewing plane is nearer than the preset near point (e.g., in the 40s, 50s, or 60s mode), the lenses adjust their diopters to decrease the wearer’s overall accommodation ability. This adjustment shifts the focal point beyond the retina, resulting in presbyopic blur. The detailed workflow can be found in Figure 5.
Figure 3. Approach and interface for presbyopia simulation using tunable lenses. (A) Illustration of the primary algorithm: The tunable lenses simulate presbyopic blur by functioning as concave lenses to reduce the overall diopter of the wearer when the viewing distance is less than a preset near point (e.g., 25 cm for 40s mode). Beyond this point, the lenses work as corrective mode adjust dynamically based on viewing distance. (B) Operational graphical interface for individual calibration, allowing separate focal power correction for each eye to accommodate wearers with refractive errors.
@@
To initialize the presbyopic simulation, tunable lenses are precisely individually adjusted to operate as concave lenses, adjusting the near point of accommodation of the wearer. This adjustment alters the overall diopter strength of the user’s vision to a preset simulation mode. In the case of a 20-year-old wearer (assuming the Diopter of amplitude of accommodation (AoA) is 9.75D) with normal vision in the mode of simulating 40s, the tunable lenses act as 0D lenses when the focused plane is far away more than 25cm, which enables the wearer to see things clearly as usual. When the distance from the focus plane is shorter than 25cm, the tunable lenses function as -5.75D concave lenses, which consequently reduces the overall amplitude of accommodation to 4.0D, the average for people in their 40s. Optically, concave-shaped tunable lenses diverge incoming light rays and shift the focal point to a location beyond the wearer’s retina, referring to Duane et al. (Duane, 1912, 1922; Purves et al., 2008) for a numerical rationale.
In contrast, in the case of simulating 40s, if the 20s wearer is nearsighted, the tunable lenses function as corrective lenses when the distance from the focused plane is greater than 25cm away from the wearer. Their corrective focal power here is stored as an offset. The distance between the tunable lens eyewear and the object viewed determines how the diopter strengths of the tunable lenses are adjusted. The varied tunable lenses’ states are illustrated in Figure 3. The distance from the viewing plane is continuously measured by a built-in Time-of-Flight sensor of the tunable lenses’ eyewear. This distance value is converted to the diopter adjustment for each lens to support the wearer with clear vision in real-time. When the distance between the focused plane and the tunable lens eyewear is shorter than 25cm, the tunable lenses now function as a power of ‘-5.8D + offset’ to shift the focal point of the wearer outside the retina to simulate presbyopia. The corresponding diopter settings of other age modes and wearer age combinations, calculated using the same principle illustrated for the 20s and 30s examples in Figure 5, are described in Table 2.
@@
Wearer age Diopter settings in presbyopic range (40s mode) Diopter settings in presbyopic range (50s mode) Diopter settings in presbyopic range (60s mode)
@@
Accommodation Speed Control: Our approach implements a non-linear accommodation speed control for the tunable lenses to mimic the accommodation speed and process observed in the aging human eye (Schaeffel et al., 1993). The algorithm uses a non-linear interpolation between current and target focal lengths. Key to this process are two variables: a time constant that adjusts the speed of accommodation, and a delta time based on the system’s refresh rate.
@@
Debounce: To ensure stable distance measurements and avoid sudden changes in lens adjustments, we incorporated a debounce algorithm. This algorithm maintains a circular buffer that holds the latest distance readings, with configurable options for the number of samples and the consensus threshold. For every new distance reading: (1) Insert the new value into the circular buffer while removing the oldest value, (2) tally the frequency of each distinct distance value in the buffer, (3) if any distance value reaches or surpasses the predefined threshold, it is deemed the stable reading, (4) if no value meets the threshold, the new distance is accepted as is.

### v2

@@ Abstract.
Presbyopia, a common age-related vision condition affecting most people as they age, often remains inadequately understood by those unaffected. To help bridge the gap between abstract accessibility knowledge and a more grounded appreciation of perceptual challenges, this study presents OpticalAging, an optical see-through simulation approach. Unlike VR-based methods, OpticalAging uses dynamically controlled tunable lenses to simulate the first-person visual perspective of presbyopia’s distance-dependent blur during real-world interaction, aiming to enhance awareness. While acknowledging critiques regarding simulation’s limitations in fully capturing lived experience, we position this tool as a complement to user-centered methods. Our user study (N = 19, 18-35 years old) provides validation: quantitative measurements show statistically significant changes in near points across three age modes (40s, 50s, 60s), while qualitative results suggest increases in self-reported understanding and awareness of perceptual challenges among participants. The integration of our tool into a design task showcases its potential applicability within age-inclusive design workflows when used critically alongside direct user engagement.
Visual Simulation, Presbyopia, Visual Experience Modulation, Programmable Vision, Optical See-through, Accessibility, Inclusive Design, Aged Vision, Assistive Technology, Design Toolkit
journalyear: 2026copyright: ccconference: The Augmented Humans International Conference 2026; March 16–19, 2026; Okinawa, Japanbooktitle: The Augmented Humans International Conference 2026 (AHs 2026), March 16–19, 2026, Okinawa, Japandoi: 10.1145/3795011.3795035isbn: 979-8-4007-2351-3/2026/03ccs: Human-centered computing Accessibility systems and toolsccs: Human-centered computing Accessibility design and evaluation methodsccs: Human-centered computing Systems and tools for interaction design
@@ 1. Introduction
Presbyopia, an age-related visual impairment, is a ubiquitous condition affecting nearly the entire aging population (34; U. Polat, C. Schor, J. Tong, A. Zomet, M. Lev, O. Yehezkel, A. Sterkin, and D. M. Levi (2012); W. Charman (2017); J. Wolffsohn, C. Leteneux-Pantais, S. Chiva-Razavi, S. Bentley, C. Johnson, A. Findley, C. Tolley, R. Arbuckle, J. Kommineni, and N. Tyagi (2020); K. R. Heys, M. Friedrich, and R. Truscott (2007); D. Atchison (2008)). Although traditional corrections such as spectacles or contact lenses offer functional solutions, they can introduce their own challenges (Katz et al., 2021; Wolffsohn et al., 2023). For instance, constantly needing to switch between different pairs of glasses can be cumbersome, multifocal lenses often require an adaptation period and may introduce peripheral distortion, and contact lenses might be uncomfortable or impractical for certain activities or environments (Waghmare and Jeria, 2022). This impact underscores a critical gap in inclusive design, which often stems from a lack of deep understanding among those unaffected by presbyopia (Sprung et al., 2022; Singer and Tusche, 2014). Addressing this understanding gap necessitates approaches that can help non-presbyopic individuals, particularly designers, appreciate the perceptual challenges associated with presbyopia, fostering broader awareness (Goertz et al., 2014). Bridging this gap requires tools that move beyond abstract design guidelines, helping designers develop a more grounded perspective than abstract knowledge alone might provide.
Advances in visual simulation technologies, including virtual reality (VR) and augmented reality (AR), have opened new avenues for simulating visual impairments. VR has shown versatility in simulating conditions and potentially enhancing understanding (1; D. Alexander, T. Nguyen, P. Keller, J. Orlosky, S. Brown, E. Wood, O. Ezenwoye, and W. Jirau-Rosaly (2020); J. Väyrynen, A. Colley, and J. Häkkilä (2016)). However, while VR offers versatile simulation platforms, its inherent disconnect from the physical world limits its utility for crucial design evaluation tasks that involve interacting with physical prototypes, assessing ergonomics directly, or judging legibility on actual product surfaces and packaging within real-world environments. Optical see-through approaches, in contrast, allow simulation directly within the user’s real-world context, enabling designers to experience simulated impairments while handling tangible mock-ups and interacting naturally with the intended use environment (Zhang et al., 2022; Krösl et al., 2020; Padmanaban et al., 2018; Langlotz et al., 2018).
@@
To validate our approach, we conducted a user study with three objectives: (1) quantitatively examine simulated presbyopia in three age modes (40s, 50s, and 60s); (2) assess reported changes in understanding and perspective-taking following interaction with the simulation; and (3) investigate the practical application of our simulation in real-world design scenarios with a professional designer and a design student. The results demonstrate that our approach significantly alters the near points of the users, enabling them to experience specific visual effects associated with presbyopia. Our system provides a perceptually relevant simulation of distance-dependent blur. Participants reported observations about daily challenges associated with these visual effects, suggesting increases in reported awareness and understanding. An exploratory case study with designer participants revealed the potential of our system to inform aspects of design processes, contributing towards more intuitive, user-centered solutions for aging populations.
This research contributes to the field of inclusive design support in several ways: (1) A novel optical see-through system using tunable lenses for first-person, real-time simulation of specific visual effects associated with presbyopia, enabling direct evaluation of real-world interactions. (2) A method intended to help bridge the gap between abstract accessibility guidelines and a more grounded appreciation of perceptual challenges relevant to inclusive design for aging populations. (3) Strong quantitative validation of the simulation’s ability to accurately shift near points, complemented by qualitative evidence suggesting potential enhancements in reported understanding and perspective-taking among non-presbyopic users when interpreted within the limitations of simulation. (4) Explores the practical applications of presbyopia simulation in real-world design scenarios, offering insight into how such tools might be integrated into professional design workflows when used critically and supplementally.
@@ 2.1. Presbyopia: Background and Treatments
2.2. Visual Impairment Simulation: From Static Filters to Optical See-Through
Simulation techniques range from low-fidelity static filters (54; K. Asada (2011); C. S. Glasses (2025)) and full-body aging suits to Virtual Reality (VR) environments (Jones et al., 2020; Krösl et al., 2020). While aging suits like the Third Age Suit (Hitchcock et al., 2001) and AGNES (Lavallière et al., 2017) effectively simulate reduced joint mobility and tactile loss, their visual components typically rely on static spectacles (e.g., yellow tints or fixed blur). Conversely, VR systems can render complex aberrations like glare and field loss (Alexander et al., 2020; Orlosky et al., 2024), but they rely on video pass-through (VST). This introduces latency and creates a disconnect from the physical world, limiting their utility for evaluating tangible prototypes where haptic feedback and focal depth are critical.
To address these limitations, researchers have turned to Optical See-Through (OST) approaches. For instance, Zhang et al. (Zhang et al., 2022) utilized OST eyewear to simulate glaucoma, emphasizing the value of an unmediated view of reality. However, prior OST works and static aging suits have not adequately addressed the dynamic, distance-dependent blur characteristic of presbyopia. Capturing this requires a real-time, mechanically adjustable optical system that goes beyond static diffusion.
Our approach specifically targets this gap to enable high-fidelity physical interaction studies. Unlike VST systems which may induce motion sickness or obscure fine surface details (e.g., paper texture), OpticalAging allows designers to assess legibility and ergonomics in their intended environment. Crucially, we situate this tool within the ethical framework proposed by disability scholars (Tigwell, 2021; Bennett and Rosner, 2019): we present simulation not as a replacement for lived expertise, but as a pragmatic “bridge” for designers operating under constraints of time and access, designed to highlight physiological barriers rather than replicate the full psychosocial experience of disability.
2.3. From Corrective Devices to Simulation Tools: A New Application for Tunable Lenses
Foundational work by researchers like Hasan et al. (Hasan et al., 2017), Li et al. (Li et al., 2006), and others (Chen et al., 2023; He et al., 2019; Liu et al., 2023; Milton et al., 2014; Jiang et al., 2021) has detailed the optical principles and control algorithms for these lenses. Building on this, applied research from Mompeán et al. (Mompeán et al., 2020) and Padmanabhan et al. (Padmanaban et al., 2018) demonstrated the efficacy of these lenses for dynamic presbyopia correction. Most recently, Agarwala et al. (Agarwala et al., 2022) evaluated liquid membrane-based systems for aberration control and feedback mechanisms. Zhang et al. (Zhang et al., 2024) demonstrated the psychophysical feasibility of using the ViXion01 eyewear to optically shift near points. Our work adopts this hardware framework but moves beyond technical feasibility to evaluate the system’s ecological validity and its utility as a design probe—areas not addressed in previous quantitative studies.
While Zhang et al. established technical feasibility, their work was limited to optical measurements in controlled settings (Zhang et al., 2024). The ecological validity of such systems, whether they produce perceptually meaningful effects during real-world tasks—and their practical utility as design probes remain unexplored. Our work addresses this gap by providing mixed-methods evaluation and an exploratory design deployment.
@@ 3. Simulating Presbyopia through Dynamic Adjustment of Diopters using Tunable Lenses
Table 1. Tunable Lenses (ViXion01) Specifications.
[…]
15.8 $\times$ 7.4mm -15D $\sim$ 15D 6.3mm 0 $\sim$ 45C 55 $\sim$ 70mm
@@
Algorithm: As presbyopia gradually progresses with age and individuals typically begin to notice symptoms in their 40s (Duane, 1922; Presbyopia, 2016), our simulation modes commence at this age, extending through the 50s and 60s, in line with previous research (Duane, 1912; Schwartz, 2013). To simulate presbyopic visual blur, the adjustable lenses alter the cumulative focal power of the wearer. The algorithm operates simply: If the tunable lens eyewear’s distance sensor identifies that the viewing plane is nearer than the preset near point (e.g., in the 40s, 50s, or 60s mode), the lenses adjust their diopters to decrease the wearer’s overall accommodation ability. This logic is the inverse of a corrective autofocus system. Whereas a corrective system would add positive diopter power to aid near-vision focus, our algorithm intentionally applies a negative diopter offset (e.g., -5.8D for a 20s user in 40s mode) precisely when the user attempts to focus up close. This computationally shifts the focal point behind the retina, simulating the physiological effect of presbyopia rather than correcting it. This results in the intended presbyopic blur. The detailed workflow can be found in Figure 4.
Figure 3. Approach and interface for presbyopia simulation using tunable lenses. (A) Illustration of the primary algorithm: The tunable lenses simulate presbyopic blur by functioning as concave lenses to reduce the overall diopter of the wearer when the viewing distance is less than a preset near point (e.g., 25 cm for 40s mode). Beyond this point, the lenses work as corrective mode to adjust dynamically based on viewing distance. (B) Operational graphical interface for individual calibration, allowing separate focal power correction for each eye to accommodate wearers with refractive errors.
@@
To initialize the presbyopic simulation, tunable lenses are precisely individually adjusted to operate as concave lenses, adjusting the near point of accommodation of the wearer. This adjustment alters the overall diopter strength of the user’s vision to a preset simulation mode. The specific diopter values used in our simulation modes (Table 2) are derived from the seminal accommodation amplitude data from Duane (Duane, 1912, 1922; Purves et al., 2008). Our algorithm operationalizes this clinical data into a dynamic, interactive system by calculating the required negative diopter adjustment needed to reduce a user’s baseline accommodation to match the target age’s average near point.
For instance, in the case of a 20-year-old wearer (assuming an amplitude of accommodation (AoA) of 9.75D) with normal vision in the 40s simulation mode, the tunable lenses act as 0D lenses when the focused plane is farther than 25cm, which enables the wearer to see things clearly as usual. When the distance from the focus plane is shorter than 25cm, the tunable lenses function as -5.75D concave lenses. This consequently reduces the overall amplitude of accommodation to 4.0D (the average for people in their 40s) and shifts the focal point to a location beyond the wearer’s retina.
In contrast, in the case of simulating 40s, if the 20s wearer is nearsighted, the tunable lenses function as corrective lenses when the distance from the focused plane is greater than 25cm away from the wearer. Their corrective focal power here is stored as an offset. The distance between the tunable lens eyewear and the object viewed determines how the diopter strengths of the tunable lenses are adjusted. The varied tunable lenses’ states are illustrated in Figure 3. The distance from the viewing plane is continuously measured by a built-in Time-of-Flight sensor of the tunable lenses’ eyewear. This distance value is converted to the diopter adjustment for each lens to support the wearer with clear vision in real-time. When the distance between the focused plane and the tunable lens eyewear is shorter than 25cm, the tunable lenses now function as a power of “-5.8D + offset” to shift the focal point of the wearer outside the retina to simulate presbyopia. The corresponding diopter settings of other age modes and wearer age combinations, calculated using the same principle illustrated for the 20s and 30s examples in Figure 4, are described in Table 2.
@@
Wearer Target Presbyopic Mode
Age 40s 50s 60s
@@
Accommodation Speed Control: Our approach implements a non-linear accommodation speed control for the tunable lenses to mimic the accommodation speed and process observed in the aging human eye (Schaeffel1993Inter‐individual). The algorithm uses a non-linear interpolation between current and target focal lengths. Key to this process are two variables: a time constant that adjusts the speed of accommodation, and a delta time based on the system’s refresh rate.
@@
Debounce: To ensure stable distance measurements and avoid sudden changes in lens adjustments, we incorporated a debounce algorithm. This algorithm maintains a circular buffer that holds the latest distance readings, with configurable options for the number of samples and the consensus threshold. For every new distance reading: (1) Insert the new value into the circular buffer while removing the oldest value, (2) calculate the frequency of each distinct distance value in the buffer, (3) if any distance value reaches or surpasses the predefined threshold, it is deemed the stable reading, (4) if no value meets the threshold, the new distance is accepted as is.